Look at both Target and Wal-Mart Stock prices. Using the dividends history for the last 5 years, calculate the growth rate for each stock. Show work.

Dividends per month
Target Wal-Mart
Month Year 0.66 0.53
AUG 19 0.64 0.53
MAY 19 0.64 0.53
FEB 19 0.64 0.53
NOV 18 0.64 0.52
AUG 18 0.64 0.52
MAY 18 0.62 0.52
FEB 18 0.62 0.52
NOV 17 0.62 0.51
AUG 17 0.62 0.51
MAY 17 0.6 0.51
NOV 16 0.6 0.5
AUG 16 0.6 0.5
MAY 16 0.56 0.5
FEB 16 0.56 0.5
NOV 15 0.56 0.49
AUG 15 0.56 0.49
MAY 15 0.52 0.49
FEB 15 0.52 0.49
Answer #1

Dividend for 2015 for Target= (0.56+0.56+0.52+0.52)=2.16

Dividend for 2019 for Target= (0.66+0.64+0.64+0.64)=2.58

Hence, growth rate for Target= (2.58/2.16)^(1/(2019-2015))-1=4.54%

Similarly for Walmart, dividend in 2015=0.49*4=1.96 and in 2019= 0.53*4=2.12

Hence, growth rate for Wal Mart= (2.12/1.96)^(1/4)-1=1.98%